Mango Fresh Company is a company engaged in the production of beverages from mangoes. There are 3 types of drinks produced by this company, namely Prince Mango, Queen Mango, and King Mango. These three types of drinks can be produced by mixing super mango and standard mango. The company assesses the quality of mangoes by giving values ​​from 1-10. The quality of the super mango is 10 while the quality of the standard mango is 5. Prince Mango has a minimum average requirement of the quality of mango used is 6, Queen Mango has a minimum average requirement of the quality of mango used is 8, while King Mango has an average requirement the minimum quality of mango used is 9. This Mango Fresh company has a stock of 5000kg of super mango and 10000kg of standard mango. Each processing of 1 kg of mango to make Prince Mango will cause variable cost of 7 ¢ and increase revenue by 20e, then processing 1 kg of mango to make Queen Mango will result in variable costs of 10e and increase revenue by 30 ¢, while for processing 1 kg of mango to make King Mango will result in variable cost of 20 ¢ and increase revenue by 50e. Formulate linear programming to maximize the profit that Mango Fresh Company can get!
Answer #1

Let Sj and Nj be the kgs of super mango and standard mango mixed to form the drink-j where j=p, q, k

Maximize Z = (20 - 7) * (Sp + Np) + (30 - 10) * (Sq + Nq) + (50 - 20) * (Sk + Nk)

Subject to,

Sp + Sq + Sk <= 5000

Np + Nq + Nk <= 10000

10 Sp + 5 Np >= 6 (Sp + Np) or, 4 Sp - 1 Np >= 0

10 Sq + 5 Nq >= 8 (Sq + Nq) or, 2 Sq - 3 Nq >= 0

10 Sk + 5 Nk >= 9 (Sk + Nk) or, 1 Sk - 4 Nk >= 0

Sj, Nj >= 0 for j=p, q, k

-----------------------------------------------------------

Excel formulation

Solution

Sp Np Sq Nq Sk Nk
Kgs. 2333 9333 0 0 2667 667
Revenue 20 20 30 30 50 50
Cost 7 7 10 10 20 20
Profit 13 13 20 20 30 30 251,667
s.t.
Capacity super mango 1 1 1 5000 <= 5000
Capacity std. mango 1 1 1 10000 <= 10000
Quality Prince 4 -1 0 >= 0
Quality Queen 2 -3 0 >= 0
Quality King 1 -4 0 >= 0